Michael received $1,000 as a graduation gift. He invested part of the money at 5% and the rest at 6%. His interest after one year is $58. How much did he invest at each rate?
Answer by checkley77(12844)

.06X+.05(1,000-X)=58
.06X+50-.05X=58
.01X=58-50
.01X=8
X=8/.01
X=$800 INVESTED @ 6%
1,000-800=$200 INVESTED @5%
PROOF:
.06*800+.05*200=58
48+10=58
58=58